A new model computer is advertised as being 6 h faster at a certain processing job than an older model. If 3 of the older models and 2 of the new models work simultaneously, they can finish the job in 1 h. How long would it take one of the new machines to finish the job alone?
my teacher gave me the answer of 3 hours and i have other problems like this that i could do but for some reason i cant set this one up. thank you :)

Answer by scott8148(6628)

each machine does some FRACTION of the job based on the individual times
___ the fractions add to one (the WHOLE job)


n = time for new machine , n+6 = time for old machine , 1 hr is combined time

(1/n) + (1/n) + [1/(n+6)] + [1/(n+6)] + [1/(n+6)] = 1

(2/n) + [3/(n+6)] = 1

2(n+6) + 3(n) = n(n+6) ___ 2n + 12 + 3n = n^2 + 6n ___ 0 = n^2 + n - 12

0 = (n+4)(n-3)

0 = n+4 ___ -4 = n ___ negative value NOT realistic

0 = n-3 ___ 3 = n